how do i set up the equation of this work problem? A rock concert brought in $432,500 on the sale of 9,500 tickets. If the tickets sold for $35 and $55 each, how ,any of each type of ticket were sold?
Answer by rothauserc(4718)

A rock concert brought in $432,500 on the sale of 9,500 tickets. If the tickets sold for $35 and $55 each, how ,any of each type of ticket were sold?
x + y = 9500
35x + 55y = 432500
solve first equation for x
x = 9500 - y
substitute for x in the second equation
35*(9500-y) + 55y = 432500
332500-35y+55y=432500
20y = 100000
y = 5000 and x = 4500,
4500 of the $35 tickets and 5000 of the $55 tickets